2. Find the general solution for the following differential equation
Answer #1

tay = -etis C333 +94 = Sec3a-ed +15y!! +94 = see? a) Auxillary carention : D² +9=0 =, D = £ 3; - g h = c, Casar +C₂ Sinar . 4,=Casa, Y = Simar W = 1% % . Edan S4+94 = -ex +15 cup = Aeat B y= Aer G=Aek Aek tqAed +9B =-ex tis lote +9B =-ea 415 10A=-1 9B =15 Arto B = up = 1 x + 4 = C